OpenStudy (anonymous):

A basketball player scores16, 22 and24 points in three games. What is the least number of points the player must score in the fourth game to average at least 20 points per game

OpenStudy (anonymous):

solve \[\frac{16+22+24+x}{4}=20\]

OpenStudy (anonymous):

first multply by 4 to get \[62+x=80\] one step and you are done

Directrix (directrix):

And, the answer is ?

OpenStudy (anonymous):

dunno, my subtraction is not so good

OpenStudy (anonymous):

18?